Let g(x) = 11/x. Find the average rate of change, A, of g between x = -1 and x = a.
A =

Answer by Alan3354(69443)

Let g(x) = 11/x. Find the average rate of change, A, of g between x = -1 and x = a.
====================
= (g(a) - g(-1))/(a+1)
= (11/a + 11)/(a+1)
= 11((a+1)/a)/(a+1)
= 11/a